OpenStudy (anonymous):

Karen makes $5 per hour for babysitting and $12 per hour giving music lessons. One week she worked a total of 18 hours and made 139$ that day. How many hours did she spend babysitting ?

OpenStudy (anonymous):

you need to use a system of equations b= babysitting m= music lessons

OpenStudy (anonymous):

we know she worked 18 hours, so.... b+m=18

OpenStudy (anonymous):

we know how much she gets paid for each, and what she made... 12 for m 5 for b so.... 139=5b+12m

OpenStudy (anonymous):

solve b+m=18 for either b or m... we'll choose b b=18-m

OpenStudy (anonymous):

substitute that in for the second equation, 139=5b+12m b=18-m 139=5(18-m)+12m

OpenStudy (anonymous):

139=90-5m+12m 49=7m m=7

OpenStudy (anonymous):

since we know m... we can plug that into the easier equation, b+m=18 b+7=18 b=11

OpenStudy (anonymous):

so, since b=11 and m=7, we know that she babysat for 11 hours and gave music lessons for 7 hours

OpenStudy (anonymous):

thanks ! i gotcha now.

OpenStudy (anonymous):

cool, good luck

OpenStudy (anonymous):

Chris put 1,500 $ in a savings account at an annual interest rate of 5%. If Chris does not deposit or withdraw any money, what is the amount of interest Chris will earn the first year her money is in the savings account ?